# Optimal control of stochastic reaction networks with entropic control cost and emergence of mode-switching strategies

### v1

@@ II.8 Controlling epidemic outbreak
Figure 4: The solution of the minimum total infection problem for stochastic SIR model for different values of $R_{0}\in\{2,4,8\}$. (a) Heat maps of the value function $V(n,1)$ plotted on the $n_{S}$$n_{I}$ plane with $\beta=1$ for varying $R_{0}$. Light color represents a large value. (b) The gradient ¯$\overline{\nabla}_{s_{1}}V(n,1)$ of the value function in the infection direction $s_{1}$, which determines the optimally controled infection rate as $k^{\dagger}_{1}(n)=k^{0}_{1}\exp(\overline{\nabla}_{s_{1}}V(n,1))$. (c) The value gradient ¯$\overline{\nabla}_{s_{2}}V(n,1)$ in the recovery direction $s_{2}$, which determines the optimally controled recovery rate as $k^{\dagger}_{2}(n)=k^{0}_{2}\exp(\overline{\nabla}_{s_{2}}V(n,1))$. (d) The value function $V(n,1)$ on the line $n_{S}=20$. (e) The gradient of the value function ¯$\overline{\nabla}_{s_{2}}V(n,1)$. Parameters are $N=50$, $\beta=1$, $k^{0}_{2}=0.01$, $k^{0}_{1}=R_{0}k^{0}_{2}$.
@@
$\underset{k(\cdot)}{\text{maximize}}\quad\beta\mathbb{E}_{k(\cdot)}\left[n_{S}(T_{\mathrm{exit}})\right]-\mathcal{D}_{KL}\left[\mathbb{P}_{k(\cdot)}\|\mathbb{P}_{k^{0}}\right].$ (42)
We numerically calculated the value function and the optimal reaction rate using Eqs. (20) and (13) for $N=50$. The results are shown in Fig. 4. The value decreases as the number of susceptible people decreases and as the number of infected people increases (Fig. 4 (a)). The optimal control of infection and recovery rates plotted in Fig. 4 (b) and (c) indicates that strong control to reduce the infection rate and to boost the recovery rate is encouraged when the number of susceptibles is large and the number of infected is small. However, if many people are already infected, strong control is no longer encouraged. Instead, almost no control over the infection and recovery rates becomes optimal, as indicated by the white regions in Fig. 4 (b) and (c). This means that almost the entire population will eventually become infected in this situation, no matter how optimally the rates are controlled, and that further investment in the control cost is not worth the potential gains.
In particular, we can observe a sharp transition between strong control and no control when $R_{0}$ is high. Figures 4 (d) and (e) show the value function $V(n,1)$ and its gradient ¯$\overline{\nabla}_{s_{2}}V(n,1)$ on the line $n_{S}=20$, which determines the optimal recovery rate $k^{\dagger}_{2}(n)=k^{0}_{2}\exp(\overline{\nabla}_{s_{2}}V(n,1))$. The value functions are approximately piecewise-linear functions $V(n,1)\approx\max\{0,a_{0}n_{S}+a_{1}n_{I}+a_{2}\}$, so the gradients and the optimal rates are approximately piecewise-constant. This transition leads to the mode switching of optimal control, which is similar to the case of the problem of maintaining diversity.
@@ III Discussion
In this paper, we proposed a new framework to formulate optimal control problems of stochastic reaction networks. We showed that the Kullback–Leibler divergence has sound properties as the control cost of RN, taming the HJB equation via its linearization through Cole-Hopf transformation. It also yields a computationally efficient expression of the optimal solutions. We demonstrated the effectiveness of our framework for three classes of biological control problems with absorbing states.
@@
There are several potential directions worth investigating. The first is risk-sensitive problems [48, 49], where not only the expectation of performance but also the variance and higher-order moments matter. For instance, as the concentration of intracellular molecules inevitably fluctuates, it is vital to suppress the fluctuation and variability when robust homeostasis is required [50]. As studied for diffusion processes [51], the current optimal control framework can be extended to incorporate risk sensitivity.
@@
Second, more realistic reaction network models can have many types or large population sizes, such as complex ecological systems and stage-structured epidemic models [52, 53, 54, 55]. The optimal control problem for large models is accompanied by high-dimensional equations. Despite the linearity of the equation, solving it is numerically challenging. The development of fast and scalable numerical algorithms is essential for addressing large-scale problems. Sampling-based techniques [24, 56] would be efficient for these high-dimensional settings.
Finally, it would be desirable if we could modify the control cost function more flexibly depending on the setting of the actual control problem. For instance, if some components of the reaction rate coefficient are constrained to lie in a certain range, or if some reactions incur a much higher control cost than the others, the control cost function has to deviate from the Kullback–Leibler divergence, resulting in the loss of the efficient computation of the optimal solutions via Cole-Hopf transformation. This limitation is analogous to the inverse proportionality condition between the weight of control cost and the noise strength, being required to solve optimal control problems for diffusion processes efficiently [24, 26]. For stochastic reaction networks, the noise strength is related to the propensity function $h_{r}(n)$, which should be used as the weight of the control cost. An iterative method with local approximation as in [57, 58] might provide a way to overcome the limitation.

### v2

@@ II.9 Controlling epidemic outbreak
Figure 6: The solution of the minimum total infection problem for stochastic SIR model for different values of $R_{0}\in\{2,4,8\}$. (a) Heat maps of the value function $V(n,1)$ plotted on the $n_{S}$$n_{I}$ plane with $\beta=1$ for varying $R_{0}$. Light color represents a large value. (b) The directional derivative $\overline{\nabla}_{s_{1}}V(n,1)$ of the value function in the infection direction $s_{1}$, which determines the optimally controled infection rate as $k^{\dagger}_{1}(n)=k^{0}_{1}\exp(\overline{\nabla}_{s_{1}}V(n,1))$. (c) The directional derivative $\overline{\nabla}_{s_{2}}V(n,1)$ of the value function in the recovery direction $s_{2}$, which determines the optimally controled recovery rate as $k^{\dagger}_{2}(n)=k^{0}_{2}\exp(\overline{\nabla}_{s_{2}}V(n,1))$. (d) The value function $V(n,1)$ on the line $n_{S}=20$. (e) The derivative of the value function $\overline{\nabla}_{s_{2}}V(n,1)$. The parameters are $N=50$, $\beta=1$, $k^{0}_{2}=0.01$, $k^{0}_{1}=R_{0}k^{0}_{2}$.
Figure 7: Comparison of the minimum total infection solution for the stochastic SIR model with different control cost functions. (a) The value function $V(n,1)$ (b,c) its derivative $\overline{\nabla}_{s_{1}}V(n,1)$ and $\overline{\nabla}_{s_{2}}V(n,1)$ with the quadratic $C_{quad}$, weighted KL $C_{wKL}$ ($w_{1}=\infty,w_{2}=1$), KL $C_{KL}$, and weighted KL $C_{wKL}$ ($w_{1}=1,w_{2}=\infty$) control cost functions. Parameters are $N=50$, $\beta=1$, $k^{0}_{2}=0.01$, $k^{0}_{1}=R_{0}k^{0}_{2}$, $R_{0}=8$.
@@
$\underset{k(\cdot)}{\text{maximize}}\quad\beta\mathbb{E}_{k(\cdot)}\left[n_{S}(T_{\mathrm{exit}})-\mathcal{C}(n(\cdot),k(\cdot))\right].$ (62)
To begin with, we assumed the KL control cost and numerically calculated the value function and the optimal reaction rate using Eqs. (37) and (32) for $N=50$. The results are shown in Fig. 6. The value decreases as the number of susceptible people decreases and as the number of infected people increases (Fig. 6 (a)). The optimal control of infection and recovery rates plotted in Fig. 6 (b) and (c) indicates that strong control to reduce the infection rate and to boost the recovery rate is encouraged when the number of susceptibles is large and the number of infected is small. However, if many people are already infected, strong control is no longer encouraged. Instead, almost no control over the infection and recovery rates becomes optimal, as indicated by the white regions in Fig. 6 (b) and (c). This means that almost the entire population will eventually become infected in this situation, no matter how optimally the rates are controlled, and that further investment in the control cost is not worth the potential gains.
In particular, we can observe a sharp transition between strong control and no control when $R_{0}$ is high. Figures 6 (d) and (e) show the value function $V(n,1)$ and its derivative $\overline{\nabla}_{s_{2}}V(n,1)$ on the line $n_{S}=20$, which determines the optimal recovery rate $k^{\dagger}_{2}(n)=k^{0}_{2}\exp(\overline{\nabla}_{s_{2}}V(n,1))$. The value functions are approximately piecewise-linear functions $V(n,1)\approx\max\{0,a_{0}n_{S}+a_{1}n_{I}+a_{2}\}$, so the derivatives and the optimal rates are approximately piecewise-constant. This transition leads to the mode switching of optimal control, which is similar to the case of the problem of maintaining diversity.
As in the previous section, we then calculated the optimal solutions with the quadratic and weighted KL control cost functions with high $R_{0}=8$. Figure 7 demonstrates the mode-switching of the optimal strategies with the quadratic and weighted KL control cost functions, as is observed with the KL control cost. Therefore, the mode switching of the optimal control is also an intrinsic feature of the minimum infection problems rather than the artifact of the KL control cost.
@@ III Discussion
In this paper, we propose a new framework for formulating optimal control problems of stochastic reaction networks. We show that the $f$-divergence possesses desirable properties as a control cost for reaction networks. In particular, when the control cost is given by the Kullback–Leibler (KL) divergence, the associated Hamilton–Jacobi–Bellman (HJB) equation can be linearized via the Cole–Hopf transformation, yielding both analytical insight and computational efficiency. We demonstrate the effectiveness of our framework through applications to three classes of biological control problems involving absorbing states.
@@
There are several potential directions worth investigating. First, our framework can be extended to risk-sensitive control problems [55, 56], where not only the expectation of performance but also its variance and higher-order moments are considered. For instance, as the concentration of intracellular molecules inevitably fluctuates, it is vital to suppress fluctuations and variability when robust homeostasis is required [57]. As studied for diffusion processes [58], risk sensitivity can be incorporated into the current framework.
@@
Second, more realistic models of reaction networks may involve many species or large population sizes, as seen in complex ecological systems and stage-structured epidemic models [59, 60, 61, 62]. The optimal control of such large-scale systems leads to high-dimensional equations. Although the HJB equation becomes linear under KL control, solving it remains computationally demanding. The development of fast and scalable numerical algorithms is therefore essential. Sampling-based techniques [28, 63] may offer efficient solutions in such high-dimensional settings.
Third, the computational advantages afforded by the Cole–Hopf transformation do not extend to general $f$-divergence control costs beyond the KL control cost. This limitation is analogous to the condition in diffusion control problems, where efficient solutions require the control cost weight to be inversely proportional to the noise intensity [28, 30]. For stochastic reaction networks, the noise strength is linked to the propensity function $h_{r}(n)$, which should be used as the weight in the control cost. As demonstrated in our examples involving the Moran and SIR models with non-KL control costs, solving the nonlinear HJB equations numerically is tractable for small-scale problems. For large-scale problems, iterative methods with local approximation, as in [64, 65], might provide a way to overcome this limitation.
Finally, the assumption of full observability of the current state does not hold in realistic scenarios. When the feedback controller does not have direct access to the true state, the optimal control problem must be reformulated to include state estimation via a filtering equation, followed by control based on the resulting belief state. Not only control but also filtering for the stochastic reaction networks is challenging due to their nonlinearity and non-Gaussian randomness [66, 67, 68]. Meanwhile, a study [69] investigated a partially observable optimal control problem for Markov jump processes with Gaussian observation noise, demonstrating that the KL control cost can be useful in deriving the optimal solution even in the partially observable setting. Developing practical algorithms for filtering and control in general stochastic reaction networks remains a promising direction for future work.